# Constraints on Axion-Like Particles from 16.5 Years of Fermi-LAT Data and Prospects for VLAST

### v1

@@ III.2 $\chi^{2}$ analysis
To investigate the signature of photon-ALP oscillation, we perform the $\chi^{2}$ analysis on the obtained SEDs of NGC 1275. Based on the 4FGL-DR4 [82], the intrinsic spectrum for the null hypothesis (labeled as w/oALPs) can be modeled as a log-parabola function:
@@
Incorporating the photon-ALP oscillation effect, the intrinsic spectrum is multiplied by the photon survival probability to obtain the oscillatory spectrum for the alternative hypothesis (labeled as wALPs):
@@
Due to the limited energy resolution of Fermi-LAT Pass 8 data at low energies ($\sim$ 10% for energies below 1 GeV), energy dispersion is introduced across the studied energy range (100 MeV500 GeV). The energy dispersion function $D_{\mathrm{eff}}\left(E^{\prime},E\right)$ is convolved with the spectral model to yield the final observed spectrum:
@@
Considering the uncertainty of the random turbulent magnetic field, we conservatively select $\chi_{\mathrm{wALPs},5}^{2}$, the 5th percentile of the best-fit $\chi_{\mathrm{wALPs}}^{2}$ values among the 100 magnetic field realization sorted in ascending order, to set constraints. This means that our limiting results are applicable to at least 95% of simulated magnetic field realizations [51].
@@ III.3 Monte Carlo Simulation
Figure 3: The results of 2500 Monte Carlo simulations based on the null hypothesis spectral model. The red vertical dashed line indicates the statistical significance threshold $\Delta\chi^{2}_{\mathrm{thr}}=9.28$ corresponding to the 95% confidence level.
@@
We simulate spectra of the null hypothesis by adopting the same energy bins as the observed SEDs. For each energy bin, the simulated flux is randomly generated from a gaussian distribution with the mean given by the best-fit intrinsic spectral model and the variance set to the observed flux error. Using this method, 2500 spectral realizations are generated. For each simulated spectrum, the aforementioned $\chi^{2}$ fitting procedure is repeated to obtain the best-fit $\chi^{2}$ values for both models (wALPs and w/oALPs). Figure 3 presents the $\chi_{\mathrm{w/oALPs}}^{2}-\chi_{\mathrm{wALPs}}^{2}$ distribution obtained from 2500 Monte Carlo simulations of the null hypothesis. The exclusion threshold is determined to be $\Delta\chi^{2}_{\mathrm{thr}}=9.28$ (red dashed line) based on 95% quantile of this null distribution. Subsequently, the ALP parameter space where $\Delta\chi^{2}$ exceeds this threshold value can be exclude at the 95% confidence level.
@@ IV.1 Fermi-LAT
Figure 4: The two-dimensional $\Delta\chi^{2}$ distribution in the parameter space of ALP masses versus photon-ALP couplings for NGC 1275. The solid black line indicates the 95% confidence level threshold ($\Delta\chi^{2}$=9.28).
Following the method introduced in Section III, we obtain the SEDs (black) of NGC 1275 based on 16.5 years of Fermi-LAT observation and make the $\chi^{2}$ fits for the wALPs (red solid line) and w/oALPs (blue dotted line) models, as shown in the Figure  2. The $\Delta\chi^{2}$ map is shown in Figure 4, where the red region indicates parameter space disfavoring the signal model (wALPs). The region with $\Delta\chi^{2}>\Delta\chi^{2}_{\mathrm{thr}}=9.28$, enclosed by the black solid line in Figure 4, has been excluded at the 95% confidence level. We find that with 16.5 years of Fermi-LAT observation, the ALP parameter space with the photon-ALP coupling of $g_{a\gamma}\gtrsim 3\times 10^{-12}\,\mathrm{GeV^{-1}}$ is excluded for the mass range $4\times 10^{-10}\,\mathrm{eV}\lesssim m_{a}\lesssim 5\times 10^{-9}\,\mathrm{eV}$, which improves than previous limits (dark green region) for 6-year data in Ref. [51] by a factor of $\sim$ 2. Additionally, as shown in Fig. 6, our result (dark orange region) excludes the previously reported “hole” region survived in Ref. [51] (dark green region). To check our method, we further repeat the above analysis with the same photon data selected in Ref. [51] and almost the same limits with this survived “hole” region are obtained. Hence, the stronger ALP limits we set are primarily due to the additional 10 years of Fermi-LAT photon data included in our study.
@@ IV.2 VLAST
Figure 5: The 68% energy resolution as a function of incident energy. The blue dashed line represents the energy resolution of Fermi-LAT (data version: P8R3_SOURCE_V3), while the red solid line shows the preliminary simulated energy resolution for VLAST. The curve in the low-energy regime corresponds to Compton scattering events, whereas the high-energy portion reflects pair-production events.
@@
Figure 6: A comparison of current experimental constraints on the parameter space of ALP. The blue-shaded regions represent exclusion limits from ground-based experiments, with darker to lighter shades corresponding to the CAST experiment [46], the ADMX experiment [38, 39, 40, 41, 42, 43, 44, 45], and the projected sensitivities of ALP-II [37] and IAXO [47], respectively. The green-shaded regions show exclusion bounds derived from astronomical observations, with darker to lighter shades representing constraints from Fermi-LAT observations of NGC 1275 [51], quasars [53], supernovae [52], and SN1987A [83], respectively. The white region denotes the parameter space where ALPs could explain the opacity of the cosmic $\gamma$-ray background [18]. In this study, exclusion bounds from observations of NGC 1275 using 16.5 years of Fermi-LAT data and 5 years of simulated VLAST data are displayed in dark and light orange, respectively.
@@
The next-generation $\gamma$-ray telescopes, such as the VLAST [63], is expected to expand the detectable the ALP parameter space by extending the energy range, improving energy resolution, and increasing the effective detection area. Compared to Fermi-LAT, VLAST represents a significant improvement in the detector performance. In the energy range above GeV, VLAST achieves approximately 5 times greater acceptance and effective area than Fermi-LAT, significantly enhancing the detection efficiency for high-energy photons. More crucially, VLAST demonstrates remarkable improvements in the energy resolution. It reaches 8% - 20% for low-energy Compton scattering events ($<$ 10 MeV) and achieves 1% - 30% for high-energy pair-conversion events (10 MeV--1 TeV) [63], representing substantial enhancement over Fermi-LAT’s 6% - 40% resolution, as seen in Figure 5. These breakthrough advancements make VLAST particularly well-suited for detecting subtle spectral features that may arise from new physics phenomena such as ALPs.
@@
We investigate the ALP detection sensitivity of the VLAST observations of NGC 1275 over a span of 5 years. Assuming the spectrum below 100 MeV still follows the log-parabola model, the photon counts for each energy bin are simulated using the best-fit intrinsic spectral model and VLAST’s exposure of 5 years. In each energy bins, the simulated counts are randomly generated from a Poisson distribution. Then we can obtain the simulated flux by ($\mathrm{count}/\mathrm{exposure}$) with flux error of ($\sqrt{\mathrm{count}}/\mathrm{exposure}$). Using this method, 100 spectral realizations are generated, and the $\chi^{2}$ fitting procedure is repeated for each.
@@
As shown in Figure 6, the light orange shaded region represents projected sensitivity of VLAST. Notably, the sensitivity of VLAST in the ALP mass range of $2\times 10^{-11}\,\mathrm{eV}\lesssim m_{a}\lesssim 1\times 10^{-7}\,\mathrm{eV}$ is expected to be stronger than that of the upcoming Axion Solar Telescope IAXO (lightskyblue region in Figure 6). Especially for $m_{a}\sim 2\times 10^{-10}\,\mathrm{eV}$, VLAST can probe the coupling down to $g_{a\gamma}\gtrsim 7\times 10^{-13}\,\mathrm{GeV^{-1}}$. In the lower mass range $m_{a}\lesssim 5\times 10^{-12}\,\mathrm{eV}$, VLAST demonstrates sensitivity to the coupling $g_{a\gamma}\gtrsim 1.5\times 10^{-11}\,\mathrm{GeV^{-1}}$. A distinct “plateau” feature appears in the parameter space for $m_{a}\lesssim 5\times 10^{-12}\,\mathrm{eV}$. As shown in Figure 1 and Equation (1), when $m_{a}\ll\omega_{\mathrm{pl}}$, the photon survival probability does not change with the ALP mass, leading to this “plateau”. Furthermore, VLAST detectable parameter space fully encompasses the region predicted by the TeV transparency phenomenon [18], indicating that future observations may provide recheck on the constraints of these theoretical predictions. The projected sensitivity of the future VLAST experiment is expected to significantly expand the detectable ALP parameter space.
@@ V Discussion
Figure 7: The photon survival probabilities predicted by different EBL models, with each model represented by distinct line colors and markers: black solid line with circles (“Franceschini” model), red solid line with triangles (“Saldaña-López” model), yellow solid line with squares (“Kneiske” model), green solid line with diamonds (“Finke”model), cyan solid line with pentagons (“Domínguez” model, the reference model used in the baseline analysis), blue solid line with crosses (“Inoue” model), and purple solid line with x-markers (“Gilmore” model).
@@
Figure 8: The photon survival probabilities under different GMF models. Left panel: Results for ALP mass significantly below the plasma frequency, showing three characteristic curves: the red solid line represents the “Jansson12” model (adopted in the baseline analysis), the green solid line corresponds to the “Pshirkov” model, and the blue solid line indicates the “UF23” model. Right panel: Results for ALP mass reaching or exceeding the plasma frequency.
@@
Figure 9: The exclusion regions in the parameter space of ALP under different GMF models. The solid lines represent the exclusion results from Fermi-LAT observations, while the colored regions indicate the detectable range of the future VLAST. Three colors are used to distinguish between different GMF models: red corresponds to the “Jansson12” model adopted in this study, green represents the “Pshirkov” model, and blue denotes the “UF23” model.
@@
The EBL in the intergalactic medium and the GMF in Milky Way, can influence the propagation of photons from extragalactic sources to Earth. In our baseline analysis, the “Dominguez” EBL model [79] and “Jansson12” GMF model  [80] are adopted. Here we conduct the comparative analysis on the potential systematic uncertainty of EBL models and GMF models.
@@
The EBL, as the remnant radiation field from galaxy formation and evolution, has been conclusively demonstrated to significantly absorb $\gamma$-ray photons through the pair-production process [79, 84]. Consequently, the influence of the EBL is specifically considered when calculating the photon survival probability. To evaluate the uncertainty of the model, comparisons are made with other EBL models [84, 85, 86, 87, 88, 89]. The differences in photon survival probability between models primarily occur at energies above 100 GeV, with a maximum discrepancy not exceeding 5% (see Figure 7). This indicates that model selection does not substantially affect our conclusions.
@@
The GMF is consisting of a large-scale regular field and a small-scale turbulent field [90, 80] and plays an important role in the photon-ALP conversion process with in the Milky Way. To assess uncertainties, two other GMF models, “Pshirkov” model [91] and the “UF23” model [92], are also considered. When the ALP mass is substantially smaller than the plasma frequency, as illustrated in the left panel of Figure 8, the photon survival probabilities for different GMF models exhibit markedly different oscillation structures within the photon energy below 100 MeV, which leads to distinct VLAST sensitivities in the low ALP mass range ($m_{a}\lesssim 1\times 10^{-11}\,\mathrm{eV}$) in Figure 9. When the ALP mass approaches or exceeds the plasma frequency in the right panel of Figure 8, the photon survival probabilities below 10 GeV vary only slightly among different GMF models. Hence, as seen in Figure 9, the uncertainties of GMF models have a negligible effect on the Fermi-LAT limits and the VLAST sensitivity in the high-mass range ($m_{a}\gtrsim 5\times 10^{-11}\,\mathrm{eV}$), but for the low-mass range ($m_{a}\lesssim 1\times 10^{-11}\,\mathrm{eV}$) the VLAST sensitivity will be influenced.
@@ VI Summary
Based on the $\gamma$-ray observations of NGC 1275, we investigate the photon-ALP oscillation structures in the Fermi-LAT measured SEDs to constrain the ALP properties and evaluate the expected detection sensitivity of VLAST over a 5-year observation period. Using 16.5 years of the Fermi-LAT observations, we exclude the parameter space with photon-ALP coupling $g_{a\gamma}\gtrsim 3\times 10^{-12}\,\mathrm{GeV^{-1}}$ within the ALP mass range of $4\times 10^{-10}\,\mathrm{eV}\lesssim m_{a}\lesssim 5\times 10^{-9}\,\mathrm{eV}$. Our results not only can exclude the previously existing “hole” survived in the 6-year observations [51], but also improves on this previous limit by a factor of 2.
By simulating the 5-year observation of NGC 1275 with VLAST, we also show that the expected sensitivity of VLAST may exceed that of the next generation Axion Solar Telescope IAXO for an ALP with the mass of $2\times 10^{-11}\,\mathrm{eV}\lesssim m_{a}\lesssim 1\times 10^{-7}\,\mathrm{eV}$. We obtain the best sensitivity of $g_{a\gamma}\sim 7\times 10^{-13}\,\mathrm{GeV^{-1}}$ at $m_{a}\sim 2\times 10^{-10}\,\mathrm{eV}$. Thanks to its detectable energy range extending below 1 MeV, VLAST will achieve sensitivity to ALP masses below $5\times 10^{-12}\,\mathrm{eV}$ with a sensitivity level of $g_{a\gamma}\gtrsim 1.5\times 10^{-11}\,\mathrm{GeV^{-1}}$ (affected by the uncertainty of the GMF model). The ALP sensitivities in this low-mass range exhibits a characteristic “plateau” feature, due to the insensitivity of photon-ALP conversion probability to the ALP mass when $m_{a}\ll\omega_{\mathrm{pl}}$. In the near future, the upcoming VLAST will extensively probe the ALP parameter space and potentially lead to important discoveries.

### v2

@@ III.2 $\chi^{2}$ analysis
To investigate the signature of photon-ALP oscillation, we perform the $\chi^{2}$ analysis on the obtained SEDs of NGC 1275. Based on the 4FGL-DR4 [92], the intrinsic spectrum for the null hypothesis (labeled as w/oALPs) can be modeled as a log-parabola function,
@@
Incorporating the photon-ALP oscillation effect, the intrinsic spectrum is multiplied by the photon survival probability to obtain the oscillatory spectrum for the alternative hypothesis (labeled as wALPs),
@@
Due to the limited energy resolution of Fermi-LAT Pass 8 data at low energies ($\sim$ 10% for energies below 1 GeV), energy dispersion is introduced across the studied energy range (100 MeV500 GeV). The energy dispersion function $D_{\mathrm{eff}}\left(E^{\prime},E\right)$ is convolved with the spectral model to yield the final observed spectrum,
@@
Considering the uncertainty of the random turbulent magnetic field, we conservatively select $\chi_{\mathrm{wALPs},5}^{2}$, the fifth percentile of the best-fit $\chi_{\mathrm{wALPs}}^{2}$ values among the 100 magnetic field realization sorted in ascending order, to set constraints. This means that our limiting results are applicable to at least 95% of simulated magnetic field realizations [58].
@@ III.3 Monte Carlo simulation
Figure 3: Results of 2500 Monte Carlo simulations based on the null hypothesis spectral model. The red vertical dashed line indicates the statistical significance threshold $\Delta\chi^{2}_{\mathrm{thr}}=9.28$ corresponding to the 95% confidence level.
@@
We simulate spectra of the null hypothesis by adopting the same energy bins as the observed SEDs. For each energy bin, the simulated flux is randomly generated from a Gaussian distribution with the mean given by the best-fit intrinsic spectral model and the variance set to the observed flux error. Using this method, 2500 spectral realizations are generated. For each simulated spectrum, the aforementioned $\chi^{2}$ fitting procedure is repeated to obtain the best-fit $\chi^{2}$ values for both models (wALPs and w/oALPs). Figure 3 presents the $\chi_{\mathrm{w/oALPs}}^{2}-\chi_{\mathrm{wALPs}}^{2}$ distribution obtained from 2500 Monte Carlo simulations of the null hypothesis. The exclusion threshold is determined to be $\Delta\chi^{2}_{\mathrm{thr}}=9.28$ (red dashed line) based on the 95% quantile of this null distribution. Subsequently, the ALP parameter space where $\Delta\chi^{2}$ exceeds this threshold value can be excluded at the 95% confidence level.
@@ IV.1 Fermi-LAT
Figure 4: Two-dimensional $\Delta\chi^{2}$ distribution in the parameter space of ALP masses versus photon-ALP couplings for NGC 1275. The solid black line indicates the 95% confidence level threshold ($\Delta\chi^{2}$=9.28).
Following the method introduced in Sec. III, we obtain the SEDs (black) of NGC 1275 based on 16.5 years of Fermi-LAT observation and make the $\chi^{2}$ fits for the wALPs (red solid line) and w/oALPs (blue dashdot line) models, as shown in Fig. 2. The $\Delta\chi^{2}$ map is shown in Fig. 4, where the red region indicates parameter space disfavoring the signal model (wALPs). The region with $\Delta\chi^{2}>\Delta\chi^{2}_{\mathrm{thr}}=9.28$, enclosed by the black solid line in Figure 4, has been excluded at the 95% confidence level. We find that with 16.5 years of Fermi-LAT observation, the ALP parameter space with the photon-ALP coupling of $g_{a\gamma}\gtrsim 3\times 10^{-12}\,\mathrm{GeV^{-1}}$ is excluded for the mass range $4\times 10^{-10}\,\mathrm{eV}\lesssim m_{a}\lesssim 5\times 10^{-9}\,\mathrm{eV}$, which improves previous limits (dark green region) for six-year data in Ref. [58] by a factor of $\sim$ 2. Additionally, as shown in Fig. 6, our result (dark orange region) excludes the previously reported “hole” region survived in Ref. [58] (dark green region). To check our method, we further repeat the above analysis with the same photon data selected in Ref. [58] and almost the same limits with this survived “hole” region are obtained. Hence, the stronger ALP limits we set are primarily due to the additional ten years of Fermi-LAT photon data included in our study.
@@ IV.2 VLAST
Figure 5: Sixty-eight percent energy resolution as a function of incident energy. The blue dashed line represents the energy resolution of Fermi-LAT (data version: P8R3_SOURCE_V3), while the red solid line shows the preliminary simulated energy resolution for VLAST. The curve in the low-energy regime corresponds to Compton scattering events, whereas the high-energy portion reflects pair-production events.
@@
Figure 6: Comparison of current experimental constraints on the parameter space of ALP. The blue-shaded regions represent exclusion limits from ground-based experiments, with darker to lighter shades corresponding to the CAST experiment [52], the ADMX experiment [44, 45, 46, 47, 48, 49, 50, 51], and the projected sensitivities of ALP-II [43] and IAXO [53], respectively. The green-shaded regions show exclusion bounds derived from astronomical observations, with darker to lighter shades representing constraints from Fermi-LAT observations of NGC 1275 [58], quasars [61], supernovae [60], and SN 1987A [59], respectively. The white region denotes the parameter space where ALPs could explain the opacity of the cosmic $\gamma$-ray background [18]. In this study, exclusion bounds from observations of NGC 1275 using 16.5 years of Fermi-LAT data and five years of simulated VLAST data are displayed in dark and light orange, respectively.
@@
The next-generation $\gamma$-ray telescopes, such as the VLAST [71], are expected to expand the detectable the ALP parameter space by extending the energy range, improving energy resolution, and increasing the effective detection area. Compared to Fermi-LAT, VLAST extends the detection energy range down to below MeV and represents a significant improvement in the instrument performance for the GeV-TeV energy range. For above 1 GeV, VLAST can achieve an effective acceptance and effective area approximately five times those of Fermi-LAT. This significantly improves the detection efficiency for high-energy photons. More crucially, VLAST demonstrates remarkable improvements in the energy resolution. It reaches 8% - 20% for low-energy Compton scattering events (<< 10 MeV) and achieves 1% - 30% for high-energy pair-conversion events (10 MeV-1 TeV) [71], representing substantial enhancement over Fermi-LAT’s 6% - 40% resolution, as seen in Fig. 5. These breakthrough advancements make VLAST particularly well suited for detecting subtle spectral features that may arise from new physics phenomena such as ALPs.
@@
We investigate the ALP detection sensitivity of the VLAST observations of NGC 1275 over a span of five years. Assuming the spectrum below 100 MeV still follows the log-parabola model, the photon counts for each energy bin are simulated using the best-fit intrinsic spectral model and VLAST’s exposure of five years. In each energy bin, the simulated counts are randomly generated from a Poisson distribution. Then we can obtain the simulated flux by ($\mathrm{count}/\mathrm{exposure}$) with flux error of ($\sqrt{\mathrm{count}}/\mathrm{exposure}$). Using this method, 100 spectral realizations are generated, and the $\chi^{2}$ fitting procedure is repeated for each.
@@
As shown in Fig. 6, the light orange shaded region represents projected sensitivity of VLAST. Notably, the sensitivity of VLAST in the ALP mass range of $2\times 10^{-11}\,\mathrm{eV}\lesssim m_{a}\lesssim 1\times 10^{-7}\,\mathrm{eV}$ is expected to be stronger than that of the upcoming Axion Solar Telescope IAXO (light sky-blue region in Fig. 6). Especially for $m_{a}\sim 2\times 10^{-10}\,\mathrm{eV}$, VLAST can probe the coupling down to $g_{a\gamma}\gtrsim 7\times 10^{-13}\,\mathrm{GeV^{-1}}$. In the lower mass range $m_{a}\lesssim 5\times 10^{-12}\,\mathrm{eV}$, VLAST demonstrates sensitivity to the coupling $g_{a\gamma}\gtrsim 1.5\times 10^{-11}\,\mathrm{GeV^{-1}}$. A distinct “plateau” feature appears in the parameter space for $m_{a}\lesssim 5\times 10^{-12}\,\mathrm{eV}$. As shown in Fig. 1 and Eq. (1), when $m_{a}\ll\omega_{\mathrm{pl}}$, the photon survival probability does not change with the ALP mass, leading to this “plateau”. Furthermore, VLAST detectable parameter space fully encompasses the region predicted by the TeV transparency phenomenon [18], indicating that future observations may provide recheck on the constraints of these theoretical predictions. The projected sensitivity of the future VLAST experiment is expected to significantly expand the detectable ALP parameter space.
@@ V Discussion
Figure 7: Photon survival probability predicted by different EBL models, with each model represented by distinct line colors and markers: black solid line with circles (“Franceschini” model), red solid line with triangles (“Saldaña-López” model), yellow solid line with squares (“Kneiske” model), green solid line with diamonds (“Finke”model), cyan solid line with pentagons (“Domínguez” model, the reference model used in the baseline analysis), blue solid line with crosses (“Inoue” model), and purple solid line with x markers (“Gilmore” model).
@@
Figure 8: Photon survival probability under different GMF models. Left panel: Results for ALP mass significantly below the plasma frequency, showing three characteristic curves: the red solid line represents the “Jansson12” model (adopted in the baseline analysis), the green solid line corresponds to the “Pshirkov” model, and the blue solid line indicates the “UF23” model. Right panel: Results for ALP mass reaching or exceeding the plasma frequency.
@@
Figure 9: Exclusion regions in the parameter space of ALP under different GMF models. The solid lines represent the exclusion results from Fermi-LAT observations, while the colored regions indicate the detectable range of the future VLAST. Three colors are used to distinguish between different GMF models: red corresponds to the “Jansson12” model adopted in this study, green represents the “Pshirkov” model, and blue denotes the “UF23” model.
@@
The EBL in the intergalactic medium and the GMF in the Milky Way, can influence the propagation of photons from extra-galactic sources to Earth. In our baseline analysis, the “Dominguez” EBL model [88] and “Jansson12” GMF model  [89] are adopted. Here we conduct the comparative analysis on the potential systematic uncertainty of EBL models and GMF models.
@@
The EBL, as the remnant radiation field from galaxy formation and evolution, has been conclusively demonstrated to significantly absorb $\gamma$-ray photons through the pair-production process [88, 93]. Consequently, the influence of the EBL is specifically considered when calculating the photon survival probability. To evaluate the uncertainty of the model, comparisons are made with other EBL models [93, 94, 95, 96, 97, 98]. The differences in photon survival probability between models primarily occur at energies above 100 GeV, with a maximum discrepancy not exceeding 5% (see Fig. 7). This indicates that model selection does not substantially affect our conclusions.
@@
The GMF is consisting of a large-scale regular field and a small-scale turbulent field [99, 89] and plays an important role in the photon-ALP conversion process within the Milky Way. To assess uncertainties, two other GMF models, the “Pshirkov” model [100] and the “UF23” model [101], are also considered. When the ALP mass is substantially smaller than the plasma frequency, as illustrated in the left panel of Fig. 8, the photon survival probability for different GMF models exhibits markedly different oscillation structures within the photon energy below 100 MeV, which leads to distinct VLAST sensitivities in the low ALP mass range ($m_{a}\lesssim 1\times 10^{-11}\,\mathrm{eV}$) in Fig. 9. When the ALP mass approaches or exceeds the plasma frequency in the right panel of Fig. 8, the photon survival probability below 10 GeV varies only slightly among different GMF models. Hence, as seen in Fig. 9, the uncertainties of GMF models have a negligible effect on the Fermi-LAT limits and the VLAST sensitivity in the high-mass range ($m_{a}\gtrsim 5\times 10^{-11}\,\mathrm{eV}$), but for the low-mass range ($m_{a}\lesssim 1\times 10^{-11}\,\mathrm{eV}$) the VLAST sensitivity will be influenced.
@@ VI Summary
Based on the $\gamma$-ray observations of NGC 1275, we investigate the photon-ALP oscillation structures in the Fermi-LAT measured SEDs to constrain the ALP properties and evaluate the expected detection sensitivity of VLAST over a five-year observation period. Using 16.5 years of the Fermi-LAT observations, we exclude the parameter space with photon-ALP coupling $g_{a\gamma}\gtrsim 3\times 10^{-12}\,\mathrm{GeV^{-1}}$ within the ALP mass range of $4\times 10^{-10}\,\mathrm{eV}\lesssim m_{a}\lesssim 5\times 10^{-9}\,\mathrm{eV}$. Our results not only can exclude the previously existing “hole” survived in the six-year observations [58], but also improve on previous limits by a factor of 2.
By simulating the five-year observation of NGC 1275 with VLAST, we also show that the expected sensitivity of VLAST may exceed that of the next generation Axion Solar Telescope IAXO for an ALP with the mass of $2\times 10^{-11}\,\mathrm{eV}\lesssim m_{a}\lesssim 1\times 10^{-7}\,\mathrm{eV}$. We obtain the best sensitivity of $g_{a\gamma}\sim 7\times 10^{-13}\,\mathrm{GeV^{-1}}$ at $m_{a}\sim 2\times 10^{-10}\,\mathrm{eV}$. Thanks to its detectable energy range extending below 1 MeV, VLAST will achieve sensitivity to ALP masses below $5\times 10^{-12}\,\mathrm{eV}$ with a sensitivity level of $g_{a\gamma}\gtrsim 1.5\times 10^{-11}\,\mathrm{GeV^{-1}}$ (affected by the uncertainty of the GMF model). The ALP sensitivities in this low-mass range exhibits a characteristic “plateau” feature, due to the insensitivity of photon-ALP conversion probability to the ALP mass when $m_{a}\ll\omega_{\mathrm{pl}}$. In the near future, the upcoming VLAST will extensively probe the ALP parameter space and potentially lead to important discoveries.